OpenStudy (anonymous):

Find the first three terms of the sequence whose nth term is n^2 +3

OpenStudy (anonymous):

1st term: \[(1)^2 + 3 = 1+3 = 4\] 2nd term\[(2)^2 + 3 = 4+3 = 7\] I'll let you try the 3rd one.

OpenStudy (anonymous):

I got it :)

OpenStudy (anonymous):

Cool.